For all values of "x" greater than 3, show how the equation: square root of x+9 equals x-3 is equivalent to x = x squared minus 6x (I do not know how to type the square root sign nor the exponent i.e. squared)
Answer by rothauserc(4718)

we are asked to show equivalence between two equations
1) (x+9)^(1/2) = x - 3
square both sides of =
x+9 = x^2 -6x +9
subtract 9 from both sides of =
x = x^2 -6x
divide both sides of = by x, we can do this because we are given that x not = 0
1 = x -6
x = 7
2) x = x^2 - 6x
divide both sides of = by x, we can do this because we are given that x not = 0
1 = x -6
x = 7
equations 1 and 2 have the same solution, they are equivalent